# Complexity of Weighted First-Order Model Counting in the Two-Variable Fragment with Counting Quantifiers: A Bound to Beat

### v1

@@ Abstract
We study the time complexity of the weighted first-order model counting (WFOMC) over the logical language with two variables and counting quantifiers. The problem is known to be solvable in time polynomial in the domain size. However, the degree of the polynomial, which turns out to be relatively high for most practical applications, has never been properly addressed. First, we formulate a time complexity bound for the existing techniques for solving WFOMC with counting quantifiers. The bound is already known to be a polynomial with its degree depending on the number of cells of the input formula. We observe that the number of cells depends, in turn, exponentially on the parameter of the counting quantifiers appearing in the formula. Second, we propose a new approach to dealing with counting quantifiers, reducing the exponential dependency to a quadratic one, therefore obtaining a tighter upper bound. It remains an open question whether the dependency of the polynomial degree on the counting quantifiers can be reduced further, thus making our new bound a bound to beat.
@@ 1 Introduction
The weighted first-order model counting (WFOMC) problem was originally proposed in the area of lifted inference as a method to perform probabilistic inference over statistical relational learning models on the lifted level (Van den Broeck et al., 2011). Its applications have ranged beyond (symbolic) probabilistic reasoning since then, including conjecturing recursive formulas in enumerative combinatorics (Barvínek et al., 2021) and discovering combinatorial integer sequences (Svatoš et al., 2023).
Regardless of the particular application context, WFOMC is also used to define a class of tractable (referred to as domain-liftable) modelling languages, i.e., languages which permit polynomial-time WFOMC computation in the domain size. The logical fragment with two variables was the first to be identified as such (Van den Broeck, 2011; Van den Broeck et al., 2014). Negative result proving that logic with three variables contains #$P_{1}$-complete counting problems followed (Beame et al., 2015), spawning many attempts to recover at least some of the expressive power provided by three and more variables, yet retaining the domain-liftability property.
@@
The work presented here has two major contributions. First, we study the domain-liftability result for $\textbf{C}^{2}$ in a greater detail, deriving a more concrete upper bound on the time complexity. Although the bound has already been proved to be polynomial in the domain size, the polynomial is yet to be explicitly derived. Moreover, as we will show, the polynomial’s degree depends exponentially on the particular counting quantifiers appearing in the input formula. Second, we propose a new approach to dealing with counting quantifiers, when computing WFOMC over $\textbf{C}^{2}$, which decreases the identified exponential dependency to a quadratic one. Thus, we further tighten the upper bound.
First, we review necessary preliminaries in Section 2. In Section 3, we derive an upper bound on the time complexity of computing WFOMC over $\textbf{C}^{2}$ for existing techniques. In Section 4, we continue by presenting a new technique for solving the problem, which improves the old upper bound considerably. To support our theoretical results, Section 5 contains several experimental scenarios. The runtime measurements therein show our improvement for several $\textbf{C}^{2}$ sentences and also sentences from one of the domain-liftable $\textbf{C}^{2}$ extensions. Finally, we review related works in Section 6 and we conclude in Section 7.
@@ 2 Background
Notation-wise, we adhere to the standard way of writing both algebraic and logical formulas. For readability purposes, we sometimes use $\cdot$ to denote multiplication and other times, as is also common, we drop the operation sign. We use boldface letters such as x to denote vectors and for any $n\in\mathds{N}$, $\left[n\right]$ denotes the set $\{1,2,\ldots,n\}$.
@@ 2.1 First-Order Logic
We work with a function-free subset of first-order logic (FOL). A particular language is defined by a finite set of variables $\mathcal{V}$, a finite set of constants (also called domain) $\Delta$  and a finite set of predicates $\mathcal{P}$. Assuming a predicate $P\in\mathcal{P}$ with arity $k$, we also write $P/k\in\mathcal{P}$. An atom has the form $P(t_{1},t_{2},\ldots,t_{k})$ where $P/k\in\mathcal{P}$ and $t_{i}\in\Delta$cup $\mathcal{V}$ are called terms. A literal is an atom or its negation. A set of formulas can be then defined inductively. Both atoms and literals are formulas. Given some formulas, more complex formulas may be formed by combining them using logical connectives, or by surrounding them with a universal ($\forall x$) or an existential ($\exists x$) quantifier where $x\in\mathcal{V}$.
A variable $x$ in a formula is called free if the formula contains no quantification over $x$, otherwise $x$ is called bound. A formula is called a sentence if it contains no free variables. A formula is called ground if it contains no variables.
We use the definition of truth (i.e., semantics) from Herbrand Logic (Hinrichs and Genesereth, 2006). The Herbrand Base (HB) of a language is the set of all ground atoms that can be constructed given the sets $\mathcal{P}$  and $\Delta$. A possible world, usually denoted by $\omega$, is any subset of HB. Atoms contained in a possible world $\omega$ are considered to be true, the rest, i.e., those contained in $\text{HB}\setminus\omega$, are considered false. The truth value of a more complex formula in a possible world is defined in the natural way. A possible world $\omega$ is a model of a formula $\varphi$ if $\varphi$ is satisfied in $\omega$.
@@ 2.1.1 FOL Fragments
We often do not work with the entire language of FOL, but rather some of its fragments. The simplest fragment that we work with is the logic with at most two variables, denoted as $\textbf{FO}^{2}$. As the name suggests, any formula from $\textbf{FO}^{2}$ contains at most two logical variables.
@@
Finally, the fragment that we pay most of our attention to is the two-variable logic with counting quantifiers, denoted by $\textbf{C}^{2}$. A counting quantifier is a generalization of the traditional existential quantifier. For a variable $x\in\mathcal{V}$, a quantifier of the form $(\exists^{=k}x)$, where $k\in\mathds{N}$, is allowed by our syntax. Satisfaction of formulas with counting quantifiers is defined naturally, in a similar manner to the satisfaction of cardinality constraints. For example, $\exists^{=k}x:\psi(x)$ is satisfied in $\omega$ if there are exactly $k$ constants $\{A_{1},A_{2},\ldots,A_{k}\}\subseteq\Delta$ such that $\forall i\in[k]:\omega\models\psi(A_{i})$. [^1] Both cardinality constraints and counting quantifiers can be defined to also allow inequality operators apart from equality. However, in techniques that we study, inequalities are handled by transforming them to equalities (Kuželka, 2021). Hence, for brevity, we opted for the simpler case with equality signs only.
Note the distinction between cardinality constraints and counting quantifiers. While the counting formula $\exists^{=k}x\;R(x)$ can be equivalently written using a cardinality constraint as $(|R|=k)$, the formula $\forall x\exists^{=k}y\;R(x,y)$ no longer permits such a simple transformation.
@@ 2.2 Weighted First-Order Model Counting
In the remainder of this text, it is often the case that we reduce one WFOMC computation (instance) to another over a different (larger) formula, possibly with fresh (not appearing in the original vocabulary) predicates. Even then, the assumed input language remains fixed in the context of the new WFOMC instance. Moreover, if we do not specify some of the weights for the new predicates, they are assumed to be unitary.
@@
(Valid Cell) A cell of a first-order formula $\varphi$ is a maximal consistent conjunction of literals formed from atoms in $\varphi$ using only a single variable. Moreover, a cell $C(x)$ of a first-order formula $\varphi(x,y)$ is valid, if $\varphi(t,t)\wedge C(t)$ is satisfiable for any constant $t\in\Delta$.
@@
Another domain-liftable language is $\textbf{FO}^{2}+\textbf{CC}$. WFOMC over $\textbf{FO}^{2}+\textbf{CC}$ is solved by repeated calls to an oracle solving WFOMC over $\textbf{FO}^{2}$. The number of required calls depends on the arities of predicates that appear in cardinality constraints. Consider $\Upsilon=\bigwedge_{i=1}^{m}\left(|R_{i}|=k_{i}\right)$ and let us denote
$\alpha(\Upsilon)=\prod_{i=1}^{m}{\left(arity(R_{i})+1\right)}.$
For an $\textbf{FO}^{2}+\textbf{CC}$ formula $\Gamma=\Phi\wedge\Upsilon$ such that $\Phi\in\textbf{FO}^{2}{}$ and $\Upsilon$ same as above, we will require $n^{\alpha}$ calls to the oracle (Kuželka, 2021).
@@
The first bound follows from the cell-based domain-liftability proof (Beame et al., 2015). [^2] The state-of-the-art algorithm for computing WFOMC over $\textbf{FO}^{2}$, i.e., FastWFOMC, improves the bound considerably in some cases (van Bremen and Kuželka, 2021a). However, as the improvements are not guranteed in the general case, we work with this bound as an effective worst case. Moreover, as we demonstrate in the experimental section, our new encoding described further in the text improves the FastWFOMC runtime for $\textbf{C}^{2}$ sentences reduced to $\textbf{FO}^{2}+\textbf{CC}$ as well. The second bound follows from Propositions $4$ and $5$ in Kuželka (2021).
It is important to note that Theorem 1 assumes all mathematical operations to take constant time. Hence, the theorem omits factors relating to bit complexity which Kuželka (2021) also addresses. However, those factors remain the same in all transformations that we consider. Therefore, for increased readability, we omit them. For a more detailed discussion on bit complexity, see the appendix.
@@ 2.3 Solving WFOMC with Counting Quantifiers
Yet another domain-liftable language is the language of $\textbf{C}^{2}$. WFOMC over $\textbf{C}^{2}$ is solved by a reduction to WFOMC over $\textbf{FO}^{2}+\textbf{CC}$ (Kuželka, 2021). The reduction consists of several steps that we review in the lemmas below. The lemmas concentrate results from other publications, hence, we make appropriate references to each of them.
First, we review a specialized Skolemization procedure for WFOMC (Van den Broeck et al., 2014) which turns an arbitrary $\textbf{FO}^{2}$ sentence into a $\textbf{UFO}^{2}$ sentence. Since all algorithms for solving WFOMC over $\textbf{FO}^{2}$ expect a universally quantified sentence as an input, this is a paramount procedure. Compared to the source publication, we present a slightly modified Skolemization procedure. The modification is due to Beame et al. (2015).
@@
Let $\Gamma=Q_{1}x_{1}Q_{2}x_{2}\ldots Q_{k}x_{k}\Phi(x_{1},\ldots,x_{k})$ be a first-order sentence in prenex normal form with each quantifier $Q_{i}$ being either $\forall$ or $\exists$ and $\Phi$ being a quantifier-free formula. Denote $j$ the first position of $\exists$. Let $\textbf{x}=(x_{1},\ldots,x_{j-1})$ and $\varphi(\textbf{x},x_{j})=Q_{j+1}x_{j+1}\ldots Q_{k}x_{k}\Phi$. Set
@@
Next, we present a technique to eliminate negation of a subformula without distributing it inside. The procedure builds on ideas from the Skolemization procedure and it was presented as Lemma 3.4 in Beame et al. (2015). It was also described as a relaxed Tseitin transform in Meert et al. (2016).
@@
Let $\neg\psi(\textbf{x})$ be a subformula of a first-order logic sentence $\Gamma$ with $k$ free variables $\textbf{x}=(x_{1},\ldots,x_{k})$. Let $C/k$ and $D/k$ be two fresh predicates with $w(C)=\overline{w}(C)=w(D)=1$ and $\overline{w}(D)=-1$. Denote $\Gamma^{\prime}$ the formula obtained from $\Gamma$ by replacing the subformula $\neg\psi(\textbf{x})$ with $C(\textbf{x})$. Let $\Upsilon=\left(\forall\textbf{x}\;C(\textbf{x})\vee D(\textbf{x})\right)\wedge\left(\forall\textbf{x}\;C(\textbf{x})\vee\psi(\textbf{x})\right)\wedge\left(\forall\textbf{x}\;D(\textbf{x})\vee\psi(\textbf{x})\right)$. Then, it holds that
@@
Finally, we move to dealing with counting quantifiers. We start with a single counting quantifier. The approach follows from Lemma 3 in Kuželka (2021).
@@
Finally, we present a case, which helps dealing with an arbitrary counting formula. It was originally presented as Lemma 4 in Kuželka (2021).
@@ 3 An Upper Bound for Existing Techniques
As we have already mentioned above, WFOMC($\varphi,n,w,\overline{w}$) for $\varphi\in\textbf{FO}^{2}+\textbf{CC}\;$ can be computed in time $\mathcal{O}(n^{p+1+\alpha})$, where $p$ is the number of valid cells of $\varphi$ (Beame et al., 2015; Kuželka, 2021) and $\alpha=\prod_{i=1}^{m}{\left(arity(R_{i})+1\right)}$ with $R_{1},R_{2},\ldots,R_{m}$ being all the predicates appearing in cardinality constraints. Let us see, how the bound increases when we try computing WFOMC over $\textbf{C}^{2}$.
3.1 A Worked Example on Removing Counting Quantifiers

### v2

@@ Abstract
We study the time complexity of weighted first-order model counting (WFOMC) over the logical language with two variables and counting quantifiers. The problem is known to be solvable in time polynomial in the domain size. However, the degree of the polynomial, which turns out to be relatively high for most practical applications, has never been properly addressed. First, we formulate a time complexity bound for the existing techniques for solving WFOMC with counting quantifiers. The bound is already known to be a polynomial with its degree depending on the number of cells of the input formula. We observe that the number of cells depends, in turn, exponentially on the parameters of the counting quantifiers appearing in the formula. Second, we propose a new approach to dealing with counting quantifiers, reducing the exponential dependency to a quadratic one, therefore obtaining a tighter upper bound. It remains an open question whether the dependency of the polynomial degree on the counting quantifiers can be reduced further, thus making our new bound a bound to beat.
@@ 1 Introduction
The weighted first-order model counting (WFOMC) problem was originally proposed in the area of lifted inference as a method to perform probabilistic inference over statistical relational learning (SRL) models on the lifted level (Van den Broeck et al., 2011). It allowed, among other things, fast learning of various SRL models (Van Haaren et al., 2015). However, its applications have ranged beyond (symbolic) probabilistic reasoning since then, including conjecturing recursive formulas in enumerative combinatorics (Barvínek et al., 2021) and discovering combinatorial integer sequences (Svatoš et al., 2023).
Regardless of the particular application context, WFOMC is also used to define a class of tractable (referred to as domain-liftable) modeling languages, i.e., languages that permit WFOMC computation in time polynomial in the domain size. The logical fragment with two variables was the first to be identified as such (Van den Broeck, 2011; Van den Broeck et al., 2014). Negative result proving that logic with three variables contains #$P_{1}$-complete counting problems followed (Beame et al., 2015), spawning many attempts to recover at least some of the expressive power provided by three and more variables, yet retaining the domain-liftability property.
@@
It follows from the domain-liftability of $\textbf{C}^{2}$ that WFOMC computation time over formulas from $\textbf{C}^{2}$ can be upper-bounded by a polynomial in the domain size. However, it turns out that the polynomial’s degree depends exponentially on the particular counting quantifiers appearing in the formula. In this paper, we propose a new approach to dealing with counting quantifiers when computing WFOMC over $\textbf{C}^{2}$, which decreases the degree’s dependency on the counting parameters from an exponential to a quadratic one, leading to a super-exponential speedup overall.
First, we review the necessary preliminaries in Section 2. In Section 3, we derive an upper bound on the time complexity of computing WFOMC over $\textbf{C}^{2}$ for existing techniques. In Section 4, we continue by presenting a new technique for solving the problem, which improves the old upper bound considerably. Section 5 contains several experimental scenarios to support our theoretical results. The runtime measurements therein show our improvement for several $\textbf{C}^{2}$ sentences and also sentences from one of the domain-liftable $\textbf{C}^{2}$ extensions. Finally, we review related works in Section 6 and conclude in Section 7.
@@ 2 Background
Notation-wise, we adhere to the standard way of writing both algebraic and logical formulas. For readability purposes, we sometimes use $\cdot$ to denote multiplication, and other times, as is also common, we drop the operation sign. We use boldface letters such as x to denote vectors and for any $n\in\mathds{N}$, $\left[n\right]$ denotes the set $\{1,2,\ldots,n\}$.
@@ 2.1 First-Order Logic
We work with a function-free subset of first-order logic (FOL). A particular language is defined by a finite set of variables $\mathcal{V}$, a finite set of constants (also called the domain) $\Delta$  and a finite set of predicates $\mathcal{P}$. Assuming a predicate $P\in\mathcal{P}$ with arity $k$, we also write $P/k\in\mathcal{P}$. An atom has the form $P(t_{1},t_{2},\ldots,t_{k})$ where $P/k\in\mathcal{P}$ and $t_{i}\in\Delta$cup $\mathcal{V}$ are called terms. A literal is an atom or its negation. A set of formulas can be then defined inductively. Both atoms and literals are formulas. Given some formulas, more complex formulas may be formed by combining them using logical connectives or by surrounding them with a universal ($\forall x$) or an existential ($\exists x$) quantifier where $x\in\mathcal{V}$.
A variable $x$ in a formula is called free if the formula contains no quantification over $x$; otherwise, $x$ is called bound. A formula is called a sentence if it contains no free variables. A formula is called ground if it contains no variables.
We use the definition of truth (i.e., semantics) from Herbrand Logic (Hinrichs and Genesereth, 2006). A language’s Herbrand Base (HB) is the set of all ground atoms that can be constructed given the sets $\mathcal{P}$  and $\Delta$. A possible world, usually denoted by $\omega$, is any subset of HB. Atoms contained in a possible world $\omega$ are considered to be true, the rest, i.e., those contained in $\text{HB}\setminus\omega$, are considered false. The truth value of a more complex formula in a possible world is defined naturally. A possible world $\omega$ is a model of a formula $\varphi$ (denoted by $\omega\models\varphi$) if $\varphi$ is satisfied in $\omega$.
@@ 2.1.1 FOL Fragments
We often do not work with the entire language of FOL but rather some of its fragments. The simplest fragment we work with is the logic with at most two variables, denoted as $\textbf{FO}^{2}$. As the name suggests, any formula from $\textbf{FO}^{2}$ contains at most two logical variables.
@@
Finally, the fragment that we pay most of our attention to is the two-variable logic with counting quantifiers, denoted by $\textbf{C}^{2}$. A counting quantifier is a generalization of the traditional existential quantifier. For a variable $x\in\mathcal{V}$, a quantifier of the form $(\exists^{=k}x)$, where $k\in\mathds{N}$, is allowed by our syntax. Satisfaction of formulas with counting quantifiers is defined naturally, similar to the satisfaction of cardinality constraints. For example, $\exists^{=k}x\;\psi(x)$ is satisfied in $\omega$ if there are exactly $k$ constants $\{A_{1},A_{2},\ldots,A_{k}\}\subseteq\Delta$ such that $\forall i\in[k]:\omega\models\psi(A_{i})$. [^1] Both CCs and counting quantifiers can also be defined to allow inequality operators. However, in the techniques that we study, inequalities are handled by transforming them to equalities (Kuželka, 2021). After such transformation, one must repeatedly solve the case with equalities only. Hence, for brevity, we present only that one case.
Note the distinction between cardinality constraints and counting quantifiers. While the counting formula $\exists^{=k}x\;R(x)$ can be equivalently written using a single cardinality constraint $(|R|=k)$, the formula $\forall x\exists^{=k}y\;R(x,y)$ no longer permits such a simple transformation.
@@ 2.2 Weighted First-Order Model Counting
In the remainder of this text, it is often the case that we reduce one WFOMC computation (instance) to another over a different (larger) formula, possibly with fresh (not appearing in the original vocabulary) predicates. Even then, the assumed input language remains fixed in the context of the new WFOMC instance. Moreover, if we do not specify some of the weights for the new predicates, they are assumed to be equal to 1.
@@
(Valid Cell) A cell of a first-order formula $\varphi$ is a maximal consistent conjunction of literals formed from atoms in $\varphi$ using only a single variable. Moreover, a cell $C(x)$ of a first-order formula $\varphi(x,y)$ is valid if and only if $\varphi(t,t)\wedge C(t)$ is satisfiable for any constant $t\in\Delta$.
@@
Another domain-liftable language is $\textbf{FO}^{2}+\textbf{CC}$. WFOMC over $\textbf{FO}^{2}+\textbf{CC}$ is solved by repeated calls to an oracle solving WFOMC over $\textbf{FO}^{2}$. The number of required calls depends on the arities of predicates that appear in cardinality constraints. Consider $\Upsilon=\bigwedge_{i=1}^{m}\left(|R_{i}|=k_{i}\right)$ and let us denote $\alpha(\Upsilon)=\sum_{i=1}^{m}{\left(arity(R_{i})+1\right)}.$ For an $\textbf{FO}^{2}+\textbf{CC}$ formula $\Gamma=\Phi\wedge\Upsilon$ such that $\Phi\in\textbf{FO}^{2}{}$ and $\Upsilon$ same as above, we will require $n^{\alpha}$ calls to the oracle (Kuželka, 2021).
@@
The first bound follows from the cell-based domain-liftability proof (Beame et al., 2015). [^2] The state-of-the-art algorithm for computing WFOMC over $\textbf{FO}^{2}$, i.e., FastWFOMC, improves the bound considerably in some cases (van Bremen and Kuželka, 2021a). However, as the improvements are not guaranteed in the general case, we work with this bound as an effective worst case. Moreover, as we demonstrate in the experimental section, our new encoding described further in the text improves the FastWFOMC runtime for $\textbf{C}^{2}$ sentences reduced to $\textbf{FO}^{2}+\textbf{CC}$ as well. The second bound follows from Propositions $4$ and $5$ in Kuželka (2021).
It is important to note that Theorem 1 assumes all mathematical operations to take constant time. Hence, the theorem omits factors relating to bit complexity, which Kuželka (2021) also addresses. However, those factors remain the same in all transformations that we consider. Therefore, we omit them for improved readability. For a more detailed discussion on bit complexity, see the appendix.
@@ 2.3 Solving WFOMC with Counting Quantifiers
Yet another domain-liftable language is the language of $\textbf{C}^{2}$. WFOMC over $\textbf{C}^{2}$ is solved by a reduction to WFOMC over $\textbf{FO}^{2}+\textbf{CC}$ (Kuželka, 2021). The reduction consists of several steps that we review in the lemmas below. The lemmas concentrate results from other publications. Hence, we make appropriate references to each of them.
First, we review a specialized Skolemization procedure for WFOMC (Van den Broeck et al., 2014), which turns an arbitrary $\textbf{FO}^{2}$ sentence into a $\textbf{UFO}^{2}$ sentence. Since all algorithms for solving WFOMC over $\textbf{FO}^{2}$ expect a universally quantified sentence as an input, this is a paramount procedure. Compared to the source publication, we present a slightly modified Skolemization procedure. The modification is due to Beame et al. (2015).
@@
Let $\Gamma=Q_{1}x_{1}Q_{2}x_{2}\ldots Q_{k}x_{k}\Phi(x_{1},\ldots,x_{k})$ be a first-order sentence in prenex normal form with each quantifier $Q_{i}$ being either $\forall$ or $\exists$ and $\Phi$ being a quantifier-free formula. Denote by $j$ the first position of $\exists$. Let $\textbf{x}=(x_{1},\ldots,x_{j-1})$ and $\varphi(\textbf{x},x_{j})=Q_{j+1}x_{j+1}\ldots Q_{k}x_{k}\Phi$. Set
@@
Next, we present a technique to eliminate negation of a subformula without distributing it inside. The procedure builds on ideas from the Skolemization procedure, and it was presented as Lemma 3.4 in Beame et al. (2015). It was also described as a relaxed Tseitin transform in Meert et al. (2016).
@@
Let $\neg\psi(\textbf{x})$ be a subformula of a first-order logic sentence $\Gamma$ with $k$ free variables $\textbf{x}=(x_{1},\ldots,x_{k})$. Let $C/k$ and $D/k$ be two fresh predicates with $w(C)=\overline{w}(C)=w(D)=1$ and $\overline{w}(D)=-1$. Denote $\Gamma^{\prime}$ the formula obtained from $\Gamma$ by replacing the subformula $\neg\psi(\textbf{x})$ with $C(\textbf{x})$. Let
$\Upsilon=\left(\forall\textbf{x}\;C(\textbf{x})\vee D(\textbf{x})\right)\wedge\left(\forall\textbf{x}\;C(\textbf{x})\vee\psi(\textbf{x})\right)\wedge\left(\forall\textbf{x}\;D(\textbf{x})\vee\psi(\textbf{x})\right).$
Then, it holds that
@@
Finally, we move to dealing with counting quantifiers. [^3] For brevity, the counting subformula in Lemmas 3, 4 and 5 contains only a single atom on a predicate $R$. That does not impede generality as the atom may represent a general subformula $\varphi$ equated to the atom using an additional universally quantified sentence, i.e., $\forall x\;R(x)\Leftrightarrow\varphi(x)$ or $\forall x\forall y\;R(x,y)\Leftrightarrow\varphi(x,y)$. We start with a single counting quantifier. The approach follows from Lemma 3 in Kuželka (2021).
@@
Finally, we present a case that helps deal with an arbitrary counting formula. It was originally presented as Lemma 4 in Kuželka (2021).
@@ 3 An Upper Bound for Existing Techniques
As we have already mentioned above, WFOMC($\varphi,n,w,\overline{w}$) for $\varphi\in\textbf{FO}^{2}+\textbf{CC}\;$ can be computed in time $\mathcal{O}(n^{p+1+\alpha})$, where $p$ is the number of valid cells of $\varphi$ (Beame et al., 2015; Kuželka, 2021) and $\alpha=\sum_{i=1}^{m}{\left(arity(R_{i})+1\right)}$ with $R_{1},R_{2},\ldots,R_{m}$ being all the predicates appearing in cardinality constraints. Let us see how the bound increases when computing WFOMC over $\textbf{C}^{2}$.
3.1 A Worked Example on Removing Counting